OpenStudy (anonymous):

-18=-14+x i need the steps. I know the answers -4

OpenStudy (lgbasallote):

solve for x??

OpenStudy (lgbasallote):

you just added 14 to both sides -18 + 14 = x -4 = x

OpenStudy (anonymous):

yes.